# VoicePilot: Harnessing LLMs as Speech Interfaces for Physically Assistive Robots

### v1

@@ 4.3. Demonstration with Community Members
After the modifications to the implementation in Version 2, we brought the system to a disability awareness event in Pittsburgh, PA, USA, where we demonstrated the speech interface to about 15 members of the community, some who had motor impairments or other types of disabilities. During the event, community members who wanted to interact with the robot provided commands to the feeding robot to have it feed a lab member either M&Ms or Cheerios cereal placed in two bowls of the Obi. We encouraged visitors to be creative with their requests to the robot. Most individuals who interacted with the robot provided only 1-2 commands and the interactions were short, lasting only a few minutes. Nevertheless, this process was helpful for evaluating the system, enabling us to identify where users were confused or struggled to use the interface, as well as hear feedback on what people liked and disliked.
@@ 5.1. Framework Iteration
Using insights from the Version 2 demonstration, we identified essential modifications and new components for our framework. First, we identified that adjustments were needed for the Robot Variables component due to GPT’s inconsistent modification of continuous variables spanning wide ranges. As an example, for the Obi system, when the LLM was provided a robot actuator speed range of 0 to 80 degrees per second and commanded to “move a little faster”, it would sometimes increase the speed by 5 degrees per second and sometimes by 20 degrees per second. We found that users would often not notice a change in speed for small increments and were confused if the speed increment was too large. This process unveiled that GPT lacked physical understanding of robot variables, but could be grounded by using small, discrete ranges for variables. For example, in the case of actuator speed for Obi, we could ground GPT’s understanding of what was “fast” or “slow” in a 0-5 discrete scale leading to speed adjustments that would be more consistent and in line with user expectations. Thus, we modified the Robot Variables component accordingly to suggest the grounding of physical variables in the prompt into small, discrete ranges with default values.
@@ 5.2. Implementation
We updated the implementation for the Obi robot using our latest framework, aiming to enhance user acceptance of the system. Our updated prompt is shown in Fig. A1. First, to adhere to the updated Robot Variables component on grounding of physical variables, we changed the continuous speed and acceleration ranges provided to GPT from [0, 80] degrees per second and [0, 250] degrees per second squared respectively to discrete [0, 5] ranges. The outputted values from GPT were linearly scaled to the original ranges before execution of the code on the robot.
In accordance with the new Feedback component, we made significant changes in the way users interacted with the system. Instead of holding down a button when they wanted to command the robot, users would now say a wakeup phrase, “Hey Obi,” to indicate that they were about to begin a command. The wakeup was implemented using the Porcupine Python API from PicoVoice (por, [n. d.]). Additionally, audio cues played by the system were added: the system would play a beep when it registered the user saying the wakeup phrase. It would additionally say “Got it, processing,” after the user finished speaking to indicate it had heard the user’s request and would say “Scooping now” or “Scraping now” whenever it was about to execute the respective action so that users would not be surprised by the robot’s movement. Lastly, the robot would announce “Ready for another command” once it had finished processing the user’s request.
Likewise, to satisfy the instructional materials component of the framework, we developed more detailed and standardized resources for learning how to use the system. We recorded an instructional video, included as Supporting File 1, explaining the purpose of the robot, how to talk to it, what to expect when using it, and what capabilities it has. The video also showed a simple demonstration of using the robot so users could see what eating with the robot looked like before trying it themselves. Additionally, we made a reference sheet, included as Supporting File 2, to summarize the capabilities of the robot. The intended purpose of this sheet was for users to have it on hand while using the robot in case they were unsure what the robot could do or how to communicate with the system. These instructional materials are intended to ease users into the new technology, with the expectation that these materials are temporary and won’t be needed as users become more comfortable with the system and its functions.
@@
Similar to the robot control functions, in accordance with the Robot Functions and Function Applications guidelines, we added the names of these new functions to the prompt, along with descriptions of their intended use cases.
@@ 5.3. Testing with Older Adults
To evaluate Version 3 of the implementation, we conducted a formal study at an independent living facility, Baptist Providence Point in the greater Pittsburgh, Pennsylvania metropolitan area, with 11 able-bodied older adults (5M, 6F) ranging in age from 72-91 (mean = 81.1, SD = 5.9). A subset of the participants are shown in Fig. 4. The ethnicity of all participants was white. Older adults were chosen in the study as they share more similarities with the target demographic as compared to the prior populations the interface was tested with due to their age. Our study, approved by Carnegie Mellon University’s Institutional Review Board, involved the researchers traveling to the independent living facility and conducting one-hour-long studies with each participant. Recruitment was conducted in person through an information session held at the independent living facility as well as over email, and informed consent was obtained from the participants for use of image in publication. Participants were asked how much experience they have with controlling a robot on a 7-point scale ranging from “1 = no experience” to “7 = expert user” and how they feel about using robots to help with everyday tasks ranging from “1 = very negatively” to “7 = very positively”. Participants had a median experience of 1 (IQR = 1, min = 1, max = 4) and a median attitude of 6 (IQR = 1, min = 4, max = 7).
@@
After the instructional session, the participant was asked to select three foods out of the following options: blueberries, yogurt, granola, pudding, pretzels, and Cheerios. As seen in Fig 2, these foods were added to bowls 0, 1, and 2 while bowl 3 is left empty. The first part of the study consisted of 1 practice task, eat a single scoop from any bowl, and 5 predefined tasks: (1) eat a single scoop from bowl 0 quickly, (2) eat a single large scoop from bowl 2, (3) scrape bowl 0 and eat a single scoop from bowl 0, (4) eat three scoops from bowl 1, (5) eat a single scoop from bowl 2 followed by single scoop from bowl 0. For each task, participants were instructed to only provide a single command. After each task, participants were asked a yes/no question, “Did the robot adequately complete the intended task?” If the participant answers no, they were given an additional attempt to complete the task, up to 3 total attempts. All surveys and questionnaires asked to participants during the study are included in the Appendix A.
@@
The next part of the study was a 7 minute open feeding session. If needed, we refilled the bowls with more food. The instructions provided to the participants were “Your goal is to finish the meal in a pleasant and efficient way. You are free to command the robot to eat the food however you like. Feel free to be creative in how you accomplish this goal.”
@@
At the end of the study, we asked participants for responses to the NASA TLX workload items on a 0-100 scale, the System Usability Scale (SUS), multiple 7-point Likert items shown in Appendix A.4, and 4 open ended questions included in Appendix A.6. We also audio and video recorded the entire study to capture participant reactions, feedback, and responses to the open-ended questions.
6. Final Framework and Validation
We present our final framework for integrating LLMs as assistive speech interfaces, outlined and detailed in Figure 3. The framework now comprises a total of nine components: Environment Description, Robot Functions, Function Applications, Code Specifications, Safety, Robot Variables, Instructional Materials, User Control Functions, and Feedback.
@@ 6. Final Framework and Validation
Figure 5. Participant reported success per task for each attempt. After both the practice task and the 5 predefined tasks, participants reply Yes/No to the question “Did the robot adequately complete the intended task?”
@@
Figure 5. Participant reported success per task for each attempt. After both the practice task and the 5 predefined tasks, participants reply Yes/No to the question “Did the robot adequately complete the intended task?”
@@
Both qualitative and quantitative data from our study with 11 participants at an independent living facility show that our framework and methodology were successful in creating a positively-received speech interface for the Obi feeding robot. As seen in Fig. 5, the majority of participants were able to complete the predefined tasks within 3 attempts, showing adequate accuracy for the LLM interface. From the responses to Likert items shown in Fig. 6, we find that participants found the interface easy to learn, answering with a median score of 6 (Agree) to L1 (Likert Item 1), “It was easy to learn how to speak to the robot”, with one participant (P10) saying, “I would imagine that most people would learn to use the system quickly.” They also found the robot easy to control through speech and felt in control when using the robot with median scores of 7 (Strongly Agree) for both items L2 (Likert Item 2) and L5 (Likert Item 5). P5 summarized this at the end of the study saying, “What I liked was it gave me a sense of control. For somebody who’s in an incapacitated situation, that would be very important… And it was easy to talk to.” Lastly, participants also gave a strong median rating of 7 (Strongly Agree) for items L4 (Likert Item 4) and L6 (Likert Item 6), related to enjoyment while using the robot and if they would use the robot and speech interface if they were unable to eat independently.
Using the NASA TLX scale, we find that participants generally reported low workload measures across all categories as seen in Fig. 7. Lastly, the average System Usability Scale score was 73.0 (SD = 18.6, min = 50.0, max = 95.0), which shows decent performance, but leaves room for improvement. The strong Likert item responses, workload measures, and SUS scores from older adults show promise for using LLMs as assistive interfaces for robot-assisted feeding and other physical tasks. Future testing with more individuals, especially those unable to feed themselves, is needed for full validation of the system.
@@ 7.2.1. Customization
Modifiers: Modifiers refer to an attempt by the user to customize the execution of commands, using words like “faster”, “slower”, “larger”, and “smaller”. For example, for the Obi robot, this might involve a desire for the robot to move at a faster speed while moving to the mouth. Throughout our study, both for the predefined tasks and the open feeding session, participants frequently issued commands to the robot accompanied by modifiers, but some expressed concerns with the system’s ability to consistently understand their modifiers.
@@
The quantitative data, shown in Fig. 5, specifically Task 1 and 2 which involved increasing robot speed and scoop size, highlights some of these challenges that participants faced with modifiers. For these tasks, 9/11 participants were successful after 3 attempts which shows room for improvement. P10 requested Obi to increase the speed during feeding, noting that “the change in speed seemed to be not great.” Similarly, P2 highlighted issues when asking for different bite sizes saying, “what I didn’t like was… [Obi was] not very accurate, especially in terms of sizes of scoops and what it was picking up each time.” Specifically, 6 participants expressed challenges with modifying the scoop size. We would like to note that scoop size challenges were primarily due to the lack of a perception system to dynamically change the depth of scoop based on the amount of food in the bowl. Our system instead used a naive approach and included a Robot Variable named “scoop depth” in the prompt that could be modified by the LLM to adjust the depth of the scoop but didn’t take into account the amount of food left in the bowl, leading to inconsistency in scoop sizes.
@@
Nevertheless, the importance of modifier customization was underscored by participant quotes that unveiled frustration and confusion when the system didn’t respond properly to their modifiers. For example, P5 expressed frustration with changing the scoop size saying, “Trying to get the amounts is a challenge”. We want to emphasize that modifiers are personalized, as they may carry varying interpretations for different users. Thus, the system should ideally adjust its execution to align with the user’s expectations. Possible approaches to achieve this include accurate sensing and perception of the environment and learning user preferences, with past work demonstrating LLMs are well suited for this task (Wu et al., 2023).
@@
Ability to process non-predefined commands: Predefined commands refer to the Robot Functions, Function Applications, and User Control Function components of our framework; for the Obi robot, these commands include scooping from a bowl, moving to the mouth, etc. On the other hand, non-predefined commands denote additional functionalities that users may want, such as mixing for the Obi robot, which are not explicitly defined and described to the LLM. Non-predefined commands are essential to an assistive interface because they allow for increased customization of robot motions in addition to modifiers, allowing the robot to more closely adhere to the user’s preferences. However, having an LLM understand and execute on non-predefined commands is challenging, especially while hard-coding robot trajectories in alignment with the Safety component of our framework. This challenge was apparent in our study, with 6 participants expressing difficulties in executing non-predefined commands, depending on the specific request and phrasing of the command. Some common non-predefined commands from the open-feeding session with Obi including mixing of foods and selection of a specific number of a food, for example, 2 blueberries in one scoop.
@@
While the vast majority of speech interfaces for assistive robots are unable to handle non-predefined commands, an advantage of LLMs is they have the ability to sequentially arrange predefined commands, which serve as ‘primitives’, to perform non-predefined commands. In the case of the Obi, during our human study, 5 participants requested the robot to mix foods, a non-predefined command. For 2 participants, GPT was successful as it consecutively ordered scooping functions from each bowl leading to a bite with multiple foods as seen in the sequence of images in Fig. 1. However, the inability to execute non-predefined commands for some participants led to confusion. For example, P1 remarked “Maybe with the mixing of the food together… [it] didn’t seem to know what mix meant.”
@@ 7.2.2. Multi-Step Instruction
Multi-Step Instruction refers to the capability of an LLM-based speech interface to execute sequential robot actions encompassed within a single user command to the robot. This guideline is closely tied to the Function Applications component of our framework in which we discuss how to prompt an LLM to understand how to combine robot functions sequentially. From our thematic analysis, we find that multi-step instructions are important to users as they can help increase system acceptance by reducing frustration and user workload. 5 participants commented on multi-step instructions, with a subset of quotes shown in Fig. 9. Some examples of multi-step instructions for the Obi robot are predefined Task 4, eating three scoops from bowl 1, and predefined Task 5, eating a single scoop from bowl 2 followed by a single scoop from bowl 0. As seen in Fig. 5, the system worked adequately for the respective tasks, with 10/11 and 11/11 participants completing the task within 3 attempts. However, there is still additional room for improvement as the system didn’t work for a subset of users during their initial attempts. P3 summarized their experience, saying “I thought it… probably has a little more difficulty with combining things” while P7 stated, “I gave it three commands, I think maybe that was too many.” Consistent functionality of multi-step instructions is crucial to consider when using LLM-based speech interfaces for assistive robots as they can increase the efficiency and comfort of users, while reducing fatigue and frustration. Multi-step instructions also enable users to conduct activities such as watching TV or conversing, while having confidence that the robot will continue the task without additional input for an extended period of time.

### v2

@@ 4.3. Demonstration with Community Members
After the modifications to the implementation in Version 2, we brought the system to a disability awareness event in Pittsburgh, PA, USA, where we demonstrated the speech interface to about 15 members of the community, some of whom had motor impairments or other types of disabilities. During the event, community members who wanted to interact with the robot provided commands to the feeding robot to have it feed a lab member either M&Ms or Cheerios cereal placed in two bowls of the Obi. We encouraged visitors to be creative with their requests to the robot. Most individuals who interacted with the robot provided only 1-2 commands and the interactions were short, lasting only a few minutes. Nevertheless, this process was helpful for evaluating the system, enabling us to identify where users were confused or struggled to use the interface, as well as hear feedback on what people liked and disliked.
@@ 5.1. Framework Iteration
Using insights from the Version 2 demonstration, we identified essential modifications and new components for our framework. First, we made adjustments to the Robot Variables component due to GPT’s inconsistent handling of continuous variables, which caused unpredictable speed changes that confused users. By grounding variables like actuator speed on a 0-5 scale, speed adjustments became more consistent and aligned with user expectations. This process allowed us to identify that while increasing the granularity of robot function and variables can enhance the user perception of robot customization, they can negatively impact the language model’s accuracy. More details are included in Appendix A.4.
@@ 5.2. Implementation
We updated the implementation for the Obi robot using our latest framework, aiming to enhance user acceptance of the system. Our updated prompt is shown in Fig. A1.
First, to adhere to the updated Robot Variables component on grounding of physical variables, we changed the continuous speed and acceleration ranges provided to GPT from [0, 80] degrees per second and [0, 250] degrees per second squared respectively to discrete [0, 5] ranges. Next, for the Feedback component, we changed how users interact with the system by replacing the need to hold down a button with a wakeup phrase, ”Hey Obi,” implemented via the Porcupine Python API from PicoVoice. We also added audio cues to inform users when the wakeup phrase is registered, when their request is being processed, and when the robot is about to execute or has completed an action. For the Instructional Materials component, we created an instructional video and a reference sheet, included as Supporting File 1 and 2, to help users understand and use the system effectively. These materials are expected to be temporary aids as users become familiar with the robot. More details can be found in Appendix A.5.
@@ 5.3. Testing with Older Adults
To evaluate Version 3 of the implementation, we conducted a formal study at an independent living facility, Baptist Providence Point in the greater Pittsburgh, Pennsylvania metropolitan area, with 11 non-disabled older adults (5M, 6F) ranging in age from 72-91 (mean = 81.1, SD = 5.9). A subset of the participants are shown in Fig. 4. The ethnicity of all participants was white. Older adults were chosen in the study as they share more similarities with the target demographic of individuals with physical impairments as compared to the prior populations the interface was tested with due to their age. Our study, approved by Carnegie Mellon University’s Institutional Review Board, involved the researchers traveling to the independent living facility and conducting one-hour-long studies with each participant. Recruitment was conducted in person through an information session held at the independent living facility as well as over email, and informed consent was obtained from the participants for use of image in publication. Participants were asked how much experience they have with controlling a robot on a 7-point scale ranging from “1 = no experience” to “7 = expert user” and how they feel about using robots to help with everyday tasks ranging from “1 = very negatively” to “7 = very positively”. Participants had a median experience of 1 (IQR = 1, min = 1, max = 4) and a median attitude of 6 (IQR = 1, min = 4, max = 7).
@@
After the instructional session, the participant was asked to select three foods out of the following options: blueberries, yogurt, granola, pudding, pretzels, and Cheerios. As seen in Fig 2, these foods were added to bowls 0, 1, and 2 while bowl 3 is left empty. The first part of the study consisted of 1 practice task, eat a single scoop from any bowl, and 5 predefined tasks: (1) eat a single scoop from bowl 0 quickly, (2) eat a single large scoop from bowl 2, (3) scrape bowl 0 and eat a single scoop from bowl 0, (4) eat three scoops from bowl 1, (5) eat a single scoop from bowl 2 followed by single scoop from bowl 0. For each task, participants were instructed to only provide a single command. After each task, participants were asked a yes/no question, “Did the robot adequately complete the intended task?” If the participant answered no, they were given an additional attempt to complete the task, up to 3 total attempts. For each task, interaction memory, the previous commands given by the user to the robot and the previous outputs by GPT, was provided to GPT for new attempts. All surveys and questionnaires asked to participants during the study are included in the Appendix B.
@@
The next part of the study was a 7 minute open feeding session. If needed, we refilled the bowls with more food. The instructions provided to the participants were “Your goal is to finish the meal in a pleasant and efficient way. You are free to command the robot to eat the food however you like. Feel free to be creative in how you accomplish this goal.” For this part of the study, interaction memory was provided to GPT for the entire session.
@@
At the end of the study, we asked participants for responses to the NASA TLX workload items on a 0-100 scale, the System Usability Scale (SUS), multiple 7-point Likert items shown in Appendix B.4, and 4 open ended questions included in Appendix B.6. We also audio and video recorded the entire study to capture participant reactions, feedback, and responses to the open-ended questions.
Both qualitative and quantitative data from our study show that our framework and methodology were successful in creating a positively-received speech interface for the Obi feeding robot. As seen in Fig. 5, the majority of participants were able to complete the predefined tasks within 3 attempts, showing adequate accuracy for the LLM interface. For reference, we include all participant commands in Appendix C. From the responses to Likert items shown in Fig. 6, we find that participants found the interface easy to learn, answering with a median score of 6 (Agree) to L1 (Likert Item 1), “It was easy to learn how to speak to the robot”, with one participant (P10) saying, “I would imagine that most people would learn to use the system quickly.” They also found the robot easy to control through speech and felt in control when using the robot with median scores of 7 (Strongly Agree) for both items L2 (Likert Item 2) and L5 (Likert Item 5). P5 summarized this at the end of the study saying, “What I liked was it gave me a sense of control. For somebody who’s in an incapacitated situation, that would be very important… And it was easy to talk to.” Lastly, participants also gave a strong median rating of 7 (Strongly Agree) for items L4 (Likert Item 4) and L6 (Likert Item 6), related to enjoyment while using the robot and if they would use the robot and speech interface if they were unable to eat independently.
Using the NASA TLX scale, we find that participants generally reported low workload measures across all categories as seen in Fig. 7. Lastly, the average System Usability Scale score was 73.0 (SD = 18.6, min = 50.0, max = 95.0), which shows decent performance, but leaves room for improvement. The strong Likert item responses, workload measures, and SUS scores from older adults show promise for using LLMs as assistive interfaces for robot-assisted feeding and other physical tasks. Future testing with more individuals, especially those unable to feed themselves, is needed for full validation of the system. To aid researchers in robot-assisted feeding, we additionally include bite-timing metrics in Appendix D.
@@
Figure 5. Participant self-reported success per task for each attempt. After both the practice task and the 5 predefined tasks, participants reply Yes/No to the question “Did the robot adequately complete the intended task?”
[…]
@@ 6. Reflection
6.1. Final Framework
We present our final framework for integrating LLMs as assistive speech interfaces, outlined and detailed in Figure 3. The framework now comprises a total of nine components: Environment Description, Robot Functions, Function Applications, Code Specifications, Safety, Robot Variables, Instructional Materials, User Control Functions, and Feedback. In the figure, we indicate if each component pertains to prompt engineering (PE), system rollout (SR), or both prompt engineering and system rollout (PS).
@@ 6.3.1. Customization
Modifiers: Modifiers refer to an attempt by the user to customize the execution of commands, using words like “faster,” “slower,” “larger,” and “smaller.” For example, for the Obi robot, this might involve a desire for the robot to move at a faster speed while moving to the mouth. Throughout our study, both for the predefined tasks and the open feeding session, participants frequently issued commands to the robot accompanied by modifiers, but some expressed concerns with the system’s ability to consistently understand their modifiers.
@@
The quantitative data, shown in Fig. 5, specifically Task 1 and 2 which involved increasing robot speed and scoop size, highlights some of these challenges that participants faced with modifiers. For Task 1, only 6 participants were successful in one attempt, which shows room for improvement. P10 requested Obi to increase the speed during feeding, noting that “the change in speed seemed to be not great.” Similarly, P2 highlighted issues when asking for different bite sizes saying, “what I didn’t like was… [Obi was] not very accurate, especially in terms of sizes of scoops and what it was picking up each time.” Specifically, 6 participants expressed challenges with modifying the scoop size. We would like to note that scoop size challenges were primarily due to the lack of a perception system to dynamically change the depth of scoop based on the amount of food in the bowl. Our system instead used a naive approach and included a Robot Variable named “scoop depth” in the prompt that could be modified by the LLM to adjust the depth of the scoop but didn’t take into account the amount of food left in the bowl, leading to inconsistency in scoop sizes.
@@
Nevertheless, the importance of modifier customization was underscored by participant quotes that unveiled frustration and confusion when the system didn’t respond properly to their modifiers. For example, P5 expressed frustration with changing the scoop size saying, “Trying to get the amounts is a challenge.” We want to emphasize that modifiers are personalized, as they may carry varying interpretations for different users. Thus, the system should ideally adjust its execution to align with the user’s expectations. Possible approaches to achieve this include accurate sensing and perception of the environment and learning user preferences, with past work demonstrating LLMs are well suited for this task (Wu et al., 2023).
@@
Ability to process non-predefined commands: Predefined commands refer to the Robot Functions, Function Applications, and User Control Function components of our framework; for the Obi robot, these commands include scooping from a bowl, moving to the mouth, etc. On the other hand, non-predefined commands denote additional functionalities that users may want, such as mixing for the Obi robot, which are not explicitly defined and described to the LLM. Non-predefined commands are essential to an assistive interface because they allow for increased customization of robot motions in addition to modifiers, allowing the robot to more closely adhere to the user’s preferences. However, having an LLM understand and execute on non-predefined commands is challenging, especially while hard-coding robot trajectories in alignment with the Safety component of our framework. This challenge was apparent in our study, with 6 participants expressing difficulties in executing non-predefined commands, depending on the specific request and phrasing of the command. Some common non-predefined commands from the open-feeding session with Obi included mixing of foods and selection of a specific quantity of a food, such as 2 blueberries in one scoop.
@@
While the vast majority of speech interfaces for assistive robots are unable to handle non-predefined commands, an advantage of LLMs is they have the ability to sequentially arrange predefined commands, which serve as ‘primitives,’ to perform non-predefined commands. In the case of the Obi, during our human study, 5 participants requested the robot to mix foods, a non-predefined command. For 2 participants, GPT was successful, as it consecutively ordered scooping functions from each bowl leading to a bite with multiple foods as seen in the sequence of images in Fig. 1. However, the inability to execute non-predefined commands for some participants led to confusion. For example, P1 remarked “Maybe with the mixing of the food together… [it] didn’t seem to know what mix meant.”
@@ 6.3.2. Multi-Step Instruction
Multi-Step Instruction refers to the capability of an LLM-based speech interface to execute sequential robot actions encompassed within a single user command to the robot. This guideline is closely tied to the Function Applications component of our framework in which we discuss how to prompt an LLM to understand how to combine robot functions sequentially. From our thematic analysis, we find that multi-step instructions are important to users as they can help increase system acceptance by reducing frustration and user workload. 5 participants commented on multi-step instructions, with a subset of quotes shown in Fig. 9. Some examples of multi-step instructions for the Obi robot are predefined Task 4, eating three scoops from bowl 1, and predefined Task 5, eating a single scoop from bowl 2 followed by a single scoop from bowl 0. As seen in Fig. 5, the system worked adequately for the respective tasks, with 9/11 and 11/11 participants completing the task within 3 attempts. However, there is still additional room for improvement as the system didn’t work for a subset of users during their initial attempts. P3 summarized their experience, saying “I thought it… probably has a little more difficulty with combining things” while P7 stated, “I gave it three commands, I think maybe that was too many.” Consistent functionality of multi-step instructions is crucial to consider when using LLM-based speech interfaces for assistive robots as they can increase the efficiency and comfort of users, while reducing fatigue and frustration. Multi-step instructions also enable users to conduct activities such as watching TV or conversing, while having confidence that the robot will continue the task without additional input for an extended period of time.